(no subject)

John Shields (jshields@ksera)
Mon, 31 May 1999 15:47:26 -0500

Shift Summary for Day shift, 05/31/99

On Shift: J. Shields
B.Hsiung
Wah

Other persons taking up Control Room space: not many-- Jason1, Jason2,
Elizabeth. A couple others slipped in and out quickly.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of 99 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7.5 hrs

Accounting of lost time:

no beam & spent some time investigating why run aborted.

========================================================================
Runs ended this shift:

Run 13238 - cosmic run, aborted
run 13239 -cosmic. still running at end of shift

Tapes brought to FCC this shift: 0

======================================================================

Accesses during shift: interlocks down/open access
(though precious few took advantage of this)

=====================================================================

Description of Shift:

For the most part, purdy gol-dern quiet for a Monday day shift-- you'd
almost think it was a national holiday or something...

08:00 trigger hung, DDD busy. Did Kfrend, restarted trigger

08:40 looked at cosmic#summary & spotted T16 has low hot rate and
ADC distrib. (sick counter?)

08:45 Crate 11 RC 9 & 10 ped. drifting by -60 counts again?!
current run 13238

13:20 YBH & JS complete gas check,

13:55 SET took regenerator out, removed 16 ns from REG A-E source,
terminated 2nd sum-out on REG top & bottom 4413s & changed both
to update mode.
Changed several 4413s for REG A-E to upd mode

14:15 SET notes rough ADC distrib. from muon run 13208 using hole triggers
(not REG*mu trig) and doing no tracking.
high gains in REG top channels believed to be cause of rate
differences between REG top and REG bot.

14:20 SET moved REG 42 to spare 1 on transmitter card.
DC off ----> -30mV

14:46 SLOW_DAQ LV alarm goes red on RR161-N1M5 LV readback for -12V
channel 1 as 0 V.
Since RR161-N1M5 is in the L1 trigger system, YBH & JL stopped
the trigger and investigated the voltage by DVM.

14:57 Taku yelled "run is aborted". Hence, cosmic run 13238 aborts.
Reason *may* be because tape drive tps2d1 needs to be cleaned.

15:00 new run 13239 begins.

15:05 Rick Coleman calls to warn us that Tevatron is doing quite well
with 11 booster batches & 4E11 intensity. John Peopls nad Steve Holms
have decided to bring in experts to fix our magnet power supply
problem.

**IMPORTANT**: IF ALL GOES WELL, WE MAY HAVE BEAM *TONIGHT*
(instead of on Wedsnesday).

personally, we'll believe it when we see it.